OpenStudy (anonymous):

x+3 x ....... + ...... = 7 5 3

OpenStudy (anonymous):

3(x +3)/5(x) = 7 3x +9 +5x = 7 8x + 9 = 7 8x = -2 divide 8 x= -2/8 = -1/4